# Modelling Non-monotone Risk Aversion and Convex Compensation in Incomplete Markets

### v1

@@ 5 Asymptotic Analysis
Suppose the settings in Theorem 2 hold. For fixed $t\in[0,T)$, $y^{*}\in(0,\infty)$,
@@
As shown in Table 1, the behaviors of $X_{t}^{*}$ provides a clear insight into the impact of market states on the risk preferences of investors. In the scenario where $\xi_{t}\rightarrow 0$, indicating an extremely favorable market state, the positive risk-seeking term $X_{t}^{R}$ drives the optimal wealth $X_{t}^{*}$ to approach infinity and the negative risk-seeking term $X_{t}^{\bar{R}}$ tends to zero. This reflects an investor’s inclination towards risk exposure supported by the optimistic prediction of the market. The loss aversion term $X_{t}^{B}$ is given by $d_{n}e^{-r(T-t)}$, signifying that even in favorable states, there is a potential of loss aversion regarding the threshold level. The term $X_{t}^{D}$ remains zero, highlighting that in highly favorable states, the impact of non-differentiable points vanishes.
Conversely, when $\xi_{t}\rightarrow\infty$, symbolizing an extremely unfavorable market state, $X_{t}^{\bar{R}}$ drives $X_{t}^{*}$ to negative infinity while $X_{t}^{R}$ remains at zero. Under such a scenario, the loss aversion term $X_{t}^{B}$ tends to $d_{0}e^{-r(T-t)}$. Similarly as above, the non-differentiable points does not affect the optimal wealth in such an extreme scenario, whereas the first threshold level $d_{0}$ results in a positive $X_{t}^{B}$.
@@ 6 Numerical Analysis
We mainly analyze the left percentage graph at $T-t=0.1$ to study the influence of linear segments $[a_{1},a_{2}],[a_{2},a_{3}],[a_{4},a_{5}]$ and non-differentiable points $a_{2}$ and $a_{4}$ on the optimal portfolio. It is evident from the analysis that non-differentiable points invariably cause a decrease in the optimal risky investment percentage. Conversely, linear sections all lead to a significant increase in the optimal portfolio. Furthermore, at all tangent points ($a_{3}$ and $a_{5}$) where the graph transitions from linearity to strictly concave segments, we always observe a pattern of decline from previously elevated levels of the risky investment.
Comparing the left figures in Figure 3 in terms of time, we see that the risky investment percentage tends to a stable “hyperbola” shape on the whole real line and an “uptick” shape on the positive wealth. To the contrast, works of Carpenter (2000), Hodder and Jackwerth (2007) and Liang, Liu, Ma and Vinoth (2024) show that the incentive compensation deduces a “peak-valley” pattern when the preference is PHARA. The common feature of PSAHARA and PHARA is that the non-concave and non-differentiable aspects are caused by an incentive contract, and as the contract approaches its maturity, the manager pursues more eagerly the high-profit incentive by increasing the risky investment, as shown in the subfigures of $T-t=0.1$.
@@ 7.1 Optimal Portfolio
$\tilde{U}\left(x\right):=\left\{\begin{aligned} &\frac{1}{0.02}\hat{U}\left(x;2,\frac{1}{0.02^{2}},0\right)+u_{0},&&x<a_{1};\\ &\frac{1}{0.02}\hat{U}^{\prime}\left(a_{1};2,\frac{1}{0.02^{2}},0\right)\left(x-a_{1}\right)\\ &\quad+\hat{U}\left(a_{1};2,\frac{1}{0.02^{2}},0\right),&&a_{1}\leqslant x<a_{2};\\ &\frac{1}{0.22}\hat{U}\left(x;2,\frac{1}{0.22^{2}},\frac{0.2B_{T}}{0.22}\right)+u_{2},&&x\geqslant a_{2}.\\ \end{aligned}\right.$
@@
By treating the composed utility (8) as a special case of the PSAHARA utility. To better illustrate the example, we set $w=0.2,v=0.02,X_{0}=1,B_{T}=X_{0}e^{\int_{0}^{T}r_{s}\,\mathrm{d}s}$ in (5), where $X_{0}$ represents the initial fund value. We let the utility parameters be $\alpha=2,\beta=1,d=0$. In this case, the concave envelope of the composed utility is shown in Figure 4.
@@ 7.2 Numerical Analysis
Similarly as in Section 6, we can plot the risky investment amounts $\boldsymbol{\pi}_{t}^{*}$ and the risky investment percentages $\frac{\boldsymbol{\pi}_{t}^{*}}{X_{t}^{*}}$.
@@ 7.3 Empirical Study
Next, we implement our optimal portfolio strategy on the U.S. stock market. The largest challenge in doing this is to estimate the value of $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ due to the volatility “smile” and “smirk” phenomenon in the market. We include the detailed procedures of estimation in Appendix B and directly show the results in this section. We conclude that no matter what estimation method we use for volatility, due to the extreme risk-taking behaviors induced by the PSAHARA portfolio, the empirical results show great volatility, which is emphasized in the Sharpe ratio analysis.
@@
Further, we use the estimation method of implied volatility in Figure 7 and of the maximum likelihood estimator (MLE) in Figure 8. The details of estimation are included in Appendix B. As shown in Figures 68, there is always a two-peak pattern in the Sharpe ratios. This means that the portfolio sometimes makes a high profit and sometimes makes a big loss. Moreover, we see that there exist more than half of the Sharpe ratios lying below 0. The reason is that the PSAHARA portfolio leads to a highly risk-taking behavior and causes a very negative daily return in many scenarios. We show an example of the wealth process in Figure 9.
@@
A theoretical interpretation of the Sharpe ratios’ two-peak pattern is the “gambling” behavior induced by the compensation scheme. Practically speaking, the SAHARA manager under the incentive scheme optimizes his objective by investing excessive amounts in the risky assets when the current wealth is below the incentive benchmark. Such behaviors incur a great volatility, resulting in the two-peak pattern shown in the figures. The volatility is so large that though achieving overall positive returns, the Sharpe ratios stay negative. Given this firm evidence, we conclude that the PSAHARA utility can generate a high return but induce a high volatility.
Acknowledgements. Y. Liu acknowledges financial support from the research startup fund at The Chinese University of Hong Kong, Shenzhen (Grant No. UDF01003336) and is partly supported by the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (Grant No. 2023B1212010001). The authors are grateful to Gongqiu Zhang and members of the research group on financial mathematics and risk management at The Chinese University of Hong Kong, Shenzhen for their useful feedback and conversations.
@@ A.1 Proof of Proposition 1
For an increasing continuous piecewise linear function $h$, we can write $h(x)=Ax+B,\;x\in\mathcal{J}$, where $A>0,B\in\mathbb{R}$, and $\mathcal{J}\subset\mathbb{R}$ is an interval segment of the domain of $h$ such that $h$ is a linear function on $\mathcal{J}$. Hence, we have
@@ A.2 Proof of Proposition 2
As each linear segment is of the SAHARA utility class, the concave envelope $U^{**}$ is a PSAHARA utility.
@@ A.3 Proof of Theorem 1
$=\left(U^{\prime}\right)^{-1}\mid_{\left(a_{k},a_{k+1}\right)}\left(y^{*}\xi_{T}\right)$
@@
(iii) If $y^{*}\xi_{T}=\gamma_{k}^{+}$ = $\gamma_{k+1}^{-}$, (i.e. $U^{**}$ is linear on $[a_{k},a_{k+1}]$), we have that $X_{T}^{*}$ can take any value in $[a_{k},a_{k+1}]$. In this case, we define the index set
Since $\xi_{T}$ is continuously distributed, we have
$\mathbb{P}\left(y^{*}\xi_{T}\in\{\gamma_{k}^{-},\gamma_{k}^{+}\}\right)\leqslant\mathbb{P}\left(y^{*}\xi_{T}\in\{\gamma_{k}^{-},\gamma_{k}^{+}\}\cup\{\gamma_{n+1}^{-}\}\right)=0.$ (45)
Moreover, $U^{**}\left(x\right)\neq U\left(x\right)$ if and only if $x\in\bigcup_{k\in\mathcal{I}}\left(a_{k},a_{k+1}\right)$. Hence, we have
$=\mathbb{E}\left[U^{**}\left(X_{T}^{*}\right)\right].$ (46)
As a result, the optimal terminal wealth is shown in Eq. (21).
(2) We observe that:
$\xi_{T}=\xi_{t}\exp\left\{-\int_{t}^{T}\left(r_{s}+\frac{1}{2}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\right)\,\mathrm{d}s-\int_{t}^{T}\boldsymbol{\theta}_{s}^{\intercal}\,\mathrm{d}\mathbf{W}_{s}\right\}.$ (47)
Define $Z_{t,T}:=\frac{\xi_{T}}{\xi_{t}}$, which is independent of $\xi_{t}$ and log-normally distributed. Since we have already obtained the optimal terminal wealth, we proceed to replicate the strategy that achieves the terminal wealth. Observe that for any self-financing trading strategy $\{Y_{t}\}_{0\leqslant t\leqslant T}$ satisfying $\mathbb{E}\left[\int_{0}^{T}Y_{s}^{2}\,\mathrm{d}s\right]<\infty$, $\{\xi_{t}Y_{t}\}_{0\leqslant t\leqslant T}$ is a $\mathcal{F}_{t}$-martingale. Hence, we can compute the optimal wealth at time $t\in[0,T)$ by the martingale representation argument:
@@ A.4 Proof of Theorem 2
From Eq. (48), we write $X_{t}^{*}$ as a function of $\xi_{t}$: $X_{t}^{*}=X_{t}(\xi_{t})$. Apply Itô’s formula to $X_{t}^{*}=X_{t}(\xi_{t})$ and obtain:
@@
Next, we discuss how to deal with the unhedgeable risk in incomplete markets. We refer to the thought experiment of fictitious completion in Karatzas, Lehoczky, Shreve and Xu (1991). We imagine that the market is completed by $(q-m)$ risky assets with the return vector $\boldsymbol{\alpha}_{t}\in\mathbb{R}^{q-m}$ and volatility matrix $\boldsymbol{\rho}_{t}\in\mathbb{R}^{(q-m)\times q}$, where the rows of $\boldsymbol{\rho}_{t}$ are orthonormal and in the kernel of $\boldsymbol{\sigma}_{t}$. Let
$\boldsymbol{\tilde{\theta}}_{\boldsymbol{\nu},t}:=\boldsymbol{\tilde{\sigma}}_{t}^{\intercal}\left(\boldsymbol{\tilde{\sigma}}_{t}\boldsymbol{\tilde{\sigma}}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\tilde{\mu}}_{t}-r_{t}\mathbf{1}_{d}\right)\\ =\boldsymbol{\theta}_{t}+\boldsymbol{\nu}_{t},$ (55)
Lemma 2.
Proposition 3.
If we assume the Novikov condition on the solutions of Eq. (13), then $\{\boldsymbol{\pi}_{t}^{*}\}_{0\leqslant t\leqslant T}$ in Theorem 2 is the optimal portfolio in the incomplete market. That is, if
$\mathbb{E}\left[\exp\left\{\frac{1}{2}\int_{0}^{T}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\,\mathrm{d}s\right\}\right]<\infty,$ (58)
for all $\{\boldsymbol{\theta}_{t}\}_{0\leqslant t\leqslant T}$ satisfying Eq. (13), then
Proof.
We only need to show that if $\{\boldsymbol{\nu}_{t}\}_{0\leqslant t\leqslant T}$ satisfies the Novikov condition, $\mathbb{E}\left[\xi_{T}^{\boldsymbol{\nu}}X_{T}^{*}\right]\leqslant x_{0}$. Define the new probability measure $\hat{\mathbb{P}}$, with
We see that if $\{\boldsymbol{\nu}_{t}\}_{0\leqslant t\leqslant T}$ satisfies the Novikov condition, $\hat{\mathbb{P}}$ is a well-defined probability measure and $\frac{\mathrm{d}\hat{\mathbb{P}}}{\mathrm{d}\mathbb{P}}$ is the Radon–Nikodym derivative. By Girsanov’s Theorem, the components of Brownian motion $\{\hat{\mathbf{W}}_{t}\}_{0\leqslant t\leqslant T}$ under $\hat{\mathbb{P}}$ is
$=\exp\left\{-\int_{0}^{T}\left(r_{s}+\frac{1}{2}\|\boldsymbol{\theta}_{s}\|_{2}^{2}\right)\,\mathrm{d}s-\int_{t}^{T}\boldsymbol{\theta}_{s}^{\intercal}\boldsymbol{\nu}_{s}\mathrm{d}s-\int_{0}^{T}\boldsymbol{\theta}_{s}^{\intercal}\,\mathrm{d}\mathbf{\hat{W}}_{s}\right\}$
It directly follows that
Lemma 2 and Proposition 3 show that if Eq. (58) holds, the optimality of $\boldsymbol{\pi}_{t}^{*}$ in Theorem 2 holds in incomplete markets.

### v2

@@ 5.1 Asymptotic Analysis
Suppose the settings in Theorem 2 hold. For fixed $t\in[0,T)$ and $y^{*}\in(0,\infty)$,
@@
As shown in Table 1, the behaviors of $X_{t}^{*}$ provide a clear insight into the impact of market states on the risk preferences of investors. In the scenario where $\xi_{t}\rightarrow 0$, indicating an extremely favorable market state, the positive risk-seeking term $X_{t}^{R}$ drives the optimal wealth $X_{t}^{*}$ to approach infinity and the negative risk-seeking term $X_{t}^{\bar{R}}$ tends to zero. This reflects an investor’s inclination towards risk exposure supported by the optimistic prediction of the market. The loss aversion term $X_{t}^{B}$ is given by $d_{n}e^{-r(T-t)}$, signifying that even in favorable states, there is a potential of loss aversion regarding the threshold level. The term $X_{t}^{D}$ remains zero, highlighting that in highly favorable states, the impact of non-differentiable points vanishes.
Conversely, when $\xi_{t}\rightarrow\infty$, symbolizing an extremely unfavorable market state, $X_{t}^{\bar{R}}$ drives $X_{t}^{*}$ to negative infinity while $X_{t}^{R}$ remains at zero. Under such a scenario, the loss aversion term $X_{t}^{B}$ tends to $d_{0}e^{-r(T-t)}$. Similarly as above, the non-differentiable points do not affect the optimal wealth in such an extreme scenario, whereas the first threshold level $d_{0}$ results in a positive $X_{t}^{B}$.
@@ 5.2 Numerical Analysis
We mainly analyze the left graph at $T-t=0.1$ to study the influence of linear segments $[a_{1},a_{2}],[a_{2},a_{3}],[a_{4},a_{5}]$ and non-differentiable points $a_{2}$ and $a_{4}$ on the optimal portfolio. It is evident from the analysis that non-differentiable points invariably cause a decrease in the optimal risky investment percentage. Conversely, linear sections all lead to a significant increase in the optimal portfolio. Furthermore, at all tangent points ($a_{3}$ and $a_{5}$) where the graph transitions from linearity to strictly concave segments, we always observe a pattern of decline from previously elevated levels of the risky investment.
Comparing the left figures in Figure 3 in terms of time, we see that the risky investment percentage tends to a stable “hyperbola” shape on the whole real line and an “uptick” shape on the positive wealth. To the contrast, works of Carpenter (2000), Hodder and Jackwerth (2007) and Liang, Liu, Ma and Vinoth (2024) show that the incentive compensation induces a “peak-valley” pattern when the preference is PHARA. The common feature of PSAHARA and PHARA is that the non-concave and non-differentiable aspects are caused by an incentive contract, and as the contract approaches its maturity, the manager pursues more eagerly the high-profit incentive by increasing the risky investment, as shown in the subfigures of $T-t=0.1$.
@@ 6.1 Optimal Portfolio
$U^{**}\left(x\right):=\left\{\begin{aligned} &\frac{1}{0.02}\hat{U}\left(x;2,\frac{1}{0.02^{2}},0\right)+u_{0},&&x<a_{1};\\ &\frac{1}{0.02}\hat{U}^{\prime}\left(a_{1};2,\frac{1}{0.02^{2}},0\right)\left(x-a_{1}\right)\\ &\quad+\hat{U}\left(a_{1};2,\frac{1}{0.02^{2}},0\right),&&a_{1}\leqslant x<a_{2};\\ &\frac{1}{0.22}\hat{U}\left(x;2,\frac{1}{0.22^{2}},\frac{0.2B_{T}}{0.22}\right)+u_{2},&&x\geqslant a_{2}.\\ \end{aligned}\right.$
@@
By treating the composed utility (8) as a special case of the PSAHARA utility, we can apply Theorems 1 and 2 to find the corresponding optimal wealth process and the optimal portfolio. To better illustrate the example, we set $w=0.2,v=0.02,X_{0}=1,B_{T}=X_{0}e^{\int_{0}^{T}r_{s}\,\mathrm{d}s}$ in (5), where $X_{0}$ represents the initial fund value. We let the utility parameters be $\alpha=2,\beta=1,d=0$. In this case, the concave envelope of the composed utility is shown in Figure 4.
@@ 6.2 Numerical Analysis
Similarly as in Section 5.2, we can plot the risky investment amounts $\boldsymbol{\pi}_{t}^{*}$ and the risky investment percentages $\frac{\boldsymbol{\pi}_{t}^{*}}{X_{t}^{*}}$.
@@ 6.3 Empirical Study
Next, we implement our optimal portfolio strategy on the U.S. stock market. The largest challenge in doing this is to estimate the value of $\{\boldsymbol{\sigma}_{t}\}_{0\leqslant t\leqslant T}$ due to the volatility “smile” and “smirk” phenomena in the market. We include the detailed procedures of estimation in Appendix B and directly show the results in this section. We conclude that no matter what estimation method we use for volatility, due to the extreme risk-taking behaviors induced by the PSAHARA portfolio, the empirical results show great volatility, which is emphasized in the Sharpe ratio analysis.
@@
Further, we use the estimation methods of implied volatility in Figure 7 and of the maximum likelihood estimator (MLE) in Figure 8. The details of estimation are included in Appendix B. As shown in Figures 68, there is always a two-peak pattern in the Sharpe ratios. This means that the portfolio sometimes makes a high profit and sometimes makes a big loss. Moreover, we see that there exist more than half of the Sharpe ratios lying below 0. The reason is that the PSAHARA portfolio leads to a highly risk-taking behavior and causes a very negative daily return in many scenarios. We show an example of the wealth process in Figure 9.
@@
A theoretical interpretation of the Sharpe ratios’ two-peak pattern is the “gambling” behavior induced by the compensation scheme. Practically speaking, the SAHARA manager under the incentive scheme optimizes his objective by investing excessive amounts in the risky assets when the current wealth is below the incentive benchmark. Such behaviors incur a great volatility, resulting in the two-peak pattern shown in the figures. The volatility is so large that though achieving overall positive returns, the Sharpe ratios stay negative. Given this evidence, we conclude that the PSAHARA utility can generate a high return but induce a high volatility.
Acknowledgements. Y. Liu acknowledges financial support from the National Natural Science Foundation of China (Grant No. 12401624), The Chinese University of Hong Kong (Shenzhen) research startup fund (Grant No. UDF01003336) and Shenzhen Science and Technology Program (Grant No. RCBS20231211090814028) and is partly supported by the Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (Grant No. 2023B1212010001). Z. Shen acknowledges financial support from the undergraduate research awards at The Chinese University of Hong Kong (Shenzhen). The authors are grateful to Gongqiu Zhang and members of the research group on financial mathematics and risk management at The Chinese University of Hong Kong, Shenzhen for their useful feedback and conversations.
@@ A.1 Proof of Proposition 1
For an increasing continuous piecewise linear function $h$, we can write $h(x)=Ax+B,\;x\in\mathcal{J}$, where $A>0,B\in\mathbb{R}$, and $\mathcal{J}\subset\mathbb{R}$ is an interval segment such that $h$ is a linear function on $\mathcal{J}$. Hence, we have
@@ A.2 Proof of Proposition 2
As each linear segment is of the SAHARA utility family, the concave envelope $U^{**}$ is a PSAHARA utility.
@@ A.3 Proof of Theorem 1
Following Section 4.2, this subsection is going to show the final step that the terminal wealth $I\left(\mathcal{Y}_{\xi}\left(x_{0}\right)\xi_{T}\right)$ is replicable, i.e., $I\left(\mathcal{Y}_{\xi}\left(x_{0}\right)\xi_{T}\right)\in\mathcal{X}^{x_{0}}$. To this end, define the process $\{V_{t}\}_{0\leqslant t\leqslant T}$:
$V_{t}:=\xi_{t}^{-1}\mathbb{E}\left[\xi_{T}I\left(\mathcal{Y}_{\xi}(x_{0})\xi_{T}\right)|\mathcal{F}_{t}\right].$ (75)
Regarding $V_{t}$ as a function of $\xi_{t}$ and applying Itô’s formula to $\xi_{t}V_{t}$, we have
$\mathrm{d}\left(\xi_{t}V_{t}\right)$ $=\left(V_{t}+\xi_{t}\frac{\partial V_{t}}{\partial\xi_{t}}\right)\mathrm{d}\xi_{t}+\frac{1}{2}\left(2\frac{\partial V_{t}}{\partial\xi_{t}}+\xi_{t}\frac{\partial^{2}V_{t}}{\partial\xi_{t}^{2}}\right)\mathrm{d}\langle\xi\rangle_{t}$ (76)
$=\left(V_{t}+\xi_{t}\frac{\partial V_{t}}{\partial\xi_{t}}\right)\left(-\xi_{t}\boldsymbol{\theta}^{\intercal}\right)\mathrm{d}\mathbf{W}_{t}+\Box\mathrm{d}t,\quad\forall t\in\left[0,T\right],$
where we do not show the drift term $\Box$ for brevity. On the other hand, for a wealth process $\{X_{t}\}_{0\leqslant t\leqslant T}$ controlled by $\{\boldsymbol{\pi}_{t}\}_{0\leqslant t\leqslant T}$, we have
$\mathrm{d}\left(\xi_{t}X_{t}\right)=\left[\xi_{t}\boldsymbol{\pi}_{t}^{\intercal}\boldsymbol{\sigma}-\xi_{t}X_{t}\boldsymbol{\theta}^{\intercal}\right]\mathrm{d}\mathbf{W}_{t}.$ (77)
Comparing the diffusion coefficients of Eqs. (76) and (77) and letting
$\boldsymbol{\pi}_{t}=-\left(\boldsymbol{\sigma}_{t}\boldsymbol{\sigma}_{t}^{\intercal}\right)^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{1}_{m}\right)\xi_{t}\frac{\partial V_{t}}{\partial\xi_{t}},$ (78)
[…]
@@
$=I\left(y^{*}\xi_{T}\right)$
@@
(iii) If $y^{*}\xi_{T}=\gamma_{k}^{+}$ = $\gamma_{k+1}^{-}$, (i.e. $U^{**}$ is linear on $[a_{k},a_{k+1}]$), we have that $X_{T}^{*}$ can take any value in $[a_{k},a_{k+1}]$. However, we know by previous discussions that
($\mathbb{P}\left(\left\{y^{*}\xi_{T}=\gamma_{k}^{+}=\gamma_{k+1}^{-}\right\}\right)=0,\;\text{for any}\;\gamma_{k}^{+},\gamma_{k+1}^{-}\in\mathbb{R}.$ (82)
Hence we can take $X_{T}^{*}$ by an arbitrary value. This finishes the proof of the first part of Theorem 1.
Now we show explicitly the wealth process that leads to terminal wealth $I\left(y^{*}\xi_{T}\right)$. Define $Z_{t,T}:=\frac{\xi_{T}}{\xi_{t}}$, which is independent of $\xi_{t}$ and log-normally distributed. Since we have already obtained the optimal terminal wealth, we proceed to replicate the strategy that achieves the terminal wealth. Observe that for any self-financing trading strategy $\{Y_{t}\}_{0\leqslant t\leqslant T}$ satisfying $\mathbb{E}\left[\int_{0}^{T}Y_{s}^{2}\,\mathrm{d}s\right]<\infty$, $\{\xi_{t}Y_{t}\}_{0\leqslant t\leqslant T}$ is an $\{\mathcal{F}_{t}\}_{0\leqslant t\leqslant T}$-martingale. Hence, we can compute the optimal wealth at time $t\in[0,T)$ by the martingale representation argument:
@@ A.4 Proof of Theorem 2
From Eq. (78), we can write $X_{t}^{*}$ as a function of $\xi_{t}$: $X_{t}^{*}=X_{t}^{*}(\xi_{t})$. Apply Itô’s formula to $X_{t}^{*}=X_{t}^{*}(\xi_{t})$ and obtain: